At Supermarket a lady bought 5 kg of rice and 7 kg of flour for Rs 54, while her friend bought 2 kg rice and 3 kg of flour  for Rs 22. Find the cost of 4 kg of rice and 4 kg of flour

Grade:

1 Answers

srikar
9 years ago
let x = cost of 1 kg rice
    y = cost of 1 kg flour
 
5x + 7y = 54 =========> 10x + 14y = 108
2x + 3y = 22 =========> 10x + 15y = 110
 
by solving y = 2 =====> cost of 1 kg flour is 2 Rs 
               x = 8 =====> cost of 1 kg rice is 8 Rs
 
=======> 4x + 4y = 40 Rs